The sequence is defined recursively. Write the first four terms.

I really need some help with this problem please.

Answer by venugopalramana(3286)

The sequence is defined recursively. Write the first four terms.
a[1]=2 a[2]=5

and a[n]=a[n-2] - 3a[n-1] for n >= 3
PUT N=3
A3=A1-3A2=2-3*5=2-15=-13
PUT N=4
A4=A2-3A3=5-3*(-13)=5+39=44

I really need some help with this problem please.